OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. l4x +3l = 9 + 2x

OpenStudy (anonymous):

x = _____ or ______

OpenStudy (anonymous):

@sirm3d

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

It's an absolute value so 4x+3 can be negative or positive. That is why there are two answers despite everything being to be first degree. 4x+3=9+2x OR -4x-3 = 9+2x

OpenStudy (anonymous):

So you need to solve: 4x+3=9+2x --->>> 2x = 6 OR -4x-3 = 9+2x ---->>>> -12 = 6x

OpenStudy (sirm3d):

each equation must satisfy \[\large 9+2x \ge 0\]

OpenStudy (anonymous):

Ahhh... I was trying to remember where the extraneous solution aspect entered in. Thanks!

OpenStudy (sirm3d):

medal for malical for showing the solution first.